# Size-structured populations with growth fluctuations: Feynman–Kac formula and decoupling

### v1

@@ Abstract
We study a size-structured population model of proliferating cells in which biomass accumulation and binary division occur at rates modulated by fluctuating internal phenotypes. We quantify how fluctuations in internal variables that influence both growth and division shape the distribution of population phenotypes. We derive conditions under which the distributions of size and internal state decouple. Under this decoupling, population‑level expectations are obtained from lineage‑level expectations by an exponential tilting given by the Feynman–Kac formula. We further characterize weaker (ensemble‑specific) versions of decoupling that hold in the lineage or the population ensemble but not both. Finally, we provide a more general interpretation of the tilted expectations in terms of the mass-weighted phenotype distribution.
@@ 1 Introduction
There is a vast literature on these questions, which can be traced back to foundational work by Euler [3]. In the mathematics literature, work on these topics falls under the umbrella of structured population models. Sharpe and Lotka were the first to study branching models with age dependence [45]. Powell later explicitly connected these models with experimental measurements of microbial population growth [40]. Many others have explored mathematical questions, such as existence and uniqueness, primarily using tools from semigroup theory [13, 51, 52]. The importance of size regulation [47] for explaining experimental measurements was only recently appreciated [31] and has motivated considerable efforts in the biophysics community [31, 37, 54, 23, 39, 18]. These studies focus separately on the dynamics of a fluctuating phenotype or on the mechanisms of cell-size regulation.
In this paper, we study a model of size-regulated growth in which the growth rate is determined by internal variables ${\boldsymbol{X}}\in\mathbb{R}^{d}$ (e.g. expression levels) that evolve independently of cell size between divisions, and may be affected at division. We further assume that $\boldsymbol{X}$} may affect the division rate, denoted $\beta$. Within this model, we explore (1) the structure of the stationary distribution of phenotypes in both the lineage and population ensemble, and (2) representations of population statistics in terms of lineage statistics, which are based on the Feynman–Kac formula. Previous work on Feynman–Kac representations has focused on models lacking the size variable [35], but a basic fact is that the division rate $\beta$ cannot be independent of size. This is most easily seen in a model where $\lambda$ is fixed (see section 4.1). In the absence of any feedback from size to the division rate, the log size will undergo a random walk, and therefore the variance in sizes will diverge [1]. Meanwhile, the dependence of $\lambda$ and $\beta$ on $\boldsymbol{X}$} generates a potentially complex correlation structure between size and the phenotype $\boldsymbol{X}$} within a population. This makes the problem of determining the distribution of $\boldsymbol{X}$} difficult, because it requires marginalization over the size distribution.
@@
In the model studied in [20], division is prescribed by a random threshold that may depend on the previous cell sizes. Our motivation was to understand more generally the relationship between the size and phenotype distributions. This is achieved by building upon the adder-sizer model in [53, 2, 24] by combining it with the fluctuating growth rate model studied in [20, 28, 48, 21].
@@
Figure 1: (A) A diagram of the model. A cell has size ($Y$) and growth ($X$) phenotypes. Division occurs at a rate that depends on both. We have omitted the explicit dependence on the initial size in this figure for simplicity. (B) A simulation of our model in the case when $X$ is an OU process, showing the evolution of $X$ (blue) and size $Y$ (red). (C) A diagram of a growing population showing the distinction between the lineage and population distributions.
@@ 1.1 Organization of this paper
$\rho_{\rm l},\,\rho_{\rm p}$ Lineage and population joint densities; under decoupling, $\rho({\boldsymbol{x}},{\boldsymbol{y}})\!=\!\nu({\boldsymbol{x}})w({\boldsymbol{y}})$
$\nu_{\rm l},\,\nu_{\rm p}$ Lineage and population growth‑phenotype densities
$w_{\rm l},\,w_{\rm p}$ Lineage and population size‑phenotype densities
$\mathbb{E}_{\rm l}$, $\mathbb{E}_{\rm p}$ expectations with respect to $\rho_{\rm l}$ and $\rho_{\rm p}$
@@ 2 Size and phenotype structured population model
Here, we describe a class of single cell growth models in which a cell’s phenotype at any instant can be decomposed into growth phenotype ${\boldsymbol{X}}(t)\in\mathbb{R}^{d}$ and size phenotype ${\boldsymbol{Y}}(t)=(Y_{c}(t),Y_{b}(t))\in\mathbb{R}^{2}$. The growth phenotype, which could represent the concentration of growth-limiting cellular components such as ribosomes or proteins involved in antibiotic resistance, evolves independently of size between cell divisions according to some Markov process as described below. The size phenotype consists of the normalized (log) current cell size and size at the most recent division. Specifically, if $M_{b}(t)$ is the size at birth of a cell at time $t$, we set ${\boldsymbol{Y}}(t)=(Y_{c}(t),Y_{b}(t))$ where $Y_{c}=\ln(M(t)/\mathbb{E}[M_{b}])$ and $Y_{b}(t)=\ln(M_{b}(t)/\mathbb{E}[M_{b}])$, and $\mathbb{E}[M_{b}]$ is the mean steady-state birth size. Meanwhile, cells accumulate size (e.g. mass) $M(t)=e^{Y_{c}(t)}$ exponentially at a rate $\lambda({\boldsymbol{X}}(t))$ and divide at a rate $\beta({\boldsymbol{X}}(t),{\boldsymbol{Y}}(t))$ yielding two newborn cells whose phenotype is sampled from a distribution $h$.
Note that the initial size is expressed as a function of time because it changes during cell division, but remains constant within a cell cycle. The motivation for keeping track of both the current size and size at birth is that many cells exhibit size-regulation mechanisms in which the growth increment between divisions depends on the initial size. To accommodate such regulatory strategies within our modeling framework, it is necessary to include both absolute cell size and initial size.
@@ 2.1 Growth phenotype dynamics
propagate expectations; that is, $g^{f}({\boldsymbol{x}},t)=\mathbb{E}[f(\tilde{{\boldsymbol{X}}}(t))|\tilde{{\boldsymbol{X}}}(0)={\boldsymbol{x}}]$ with $g^{f}({\boldsymbol{x}},0)=f({\boldsymbol{x}})$. We use the convention that $\mathcal{L}$ (without the adjoint notation) propagates the probability density because this object will appear more often in our presentation. As explained later, we will allow the possibility that $\mathcal{L}=0$, in which case $\tilde{{\boldsymbol{X}}}(t)=\tilde{{\boldsymbol{X}}}(0)$.
@@ 2.2 Size and division dynamics
where $t_{k}$ is the $k$th division event, $({\boldsymbol{X}}(t_{k}^{-}),Y_{c}(t_{k}^{-}))$ are the growth and size phenotypes immediately before division and ${\boldsymbol{x}},{\boldsymbol{x}}+d{\boldsymbol{x}})$ is shorthand for the rectangle $x_{1},x_{1}+dx_{1})\times\cdots\times[x_{d},x_{d}+dx_{d})$.
@@
  • A linear operator $\mathcal{L}$ which is the generator of a continuous time Markov process (including the trivial case $\mathcal{L}=0$).
@@
  • A transition kernel $h:(\mathbb{R}^{d}\times\mathbb{R})\times(\mathbb{R}^{d}\times\mathbb{R})\to\mathbb{R}_{\geq 0}$ gives the phenotypes of a cell following a division event. $h(\cdot,\cdot|{\boldsymbol{x}},y_{c})$ is a probability density which respects the conservation of size at division– that is, the sum of the two newborn daughter cell sizes must equal the mother cell size immediately before division. Said another way, if $y_{c}^{\prime}$ is the log of the size of the mother cell before division and $y_{c}$ of a daughter cell, then $e^{y_{c}^{\prime}}-e^{y_{c}}$ (the mass of the other daughter cell) is equal in distribution to $e^{y_{c}}$.
@@ 2.3 Lineage distribution
The lineage distribution is obtained when we consider the distribution of phenotypes over a sequence of cells, where at each division, one of the daughter cells is discarded. The evolution of the joint density $\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}})$ of growth and size phenotypes, $({\boldsymbol{X}}_{t},{\boldsymbol{Y}}_{t})=({\boldsymbol{x}},{\boldsymbol{y}})$ along a lineage obeys the PDE
$\begin{split}\partial_{t}\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}},t)&=\mathcal{A}\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}},t)\end{split}$ (6)
@@
The boundary conditions are derived from examining the flux in the $y_{c}$ direction across the boundary at zero added size ($y_{c}=y_{b}$). Assuming $\beta({\boldsymbol{x}},{\boldsymbol{y}})$ does not diverge at $y_{c}=y_{b}$ (cells do not divide at the instant of division), we have
$\begin{split}&\lambda({\boldsymbol{x}})\rho_{\rm l}^{b}({\boldsymbol{x}},y_{b},t)=\int_{y_{b}}^{\infty}\int_{-\infty}^{y_{c}^{\prime}}\int_{\mathbb{R}^{d}}h({\boldsymbol{x}},y_{b}|{\boldsymbol{x}}^{\prime},y_{c}^{\prime})\beta({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime})\rho_{\rm l}({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime},t)d{\boldsymbol{x}}^{\prime}dy_{b}^{\prime}dy_{c}^{\prime}\end{split}$ (8)
where
$\rho_{\rm l}^{b}({\boldsymbol{x}},y_{b},t)=\rho_{\rm l}({\boldsymbol{x}},(y_{b},y_{b}),t).$ (9)
Throughout this paper, we will use the superscript $b$ to denote the distribution over newborn cells.
We are solely interested in the steady-state distributions
$\rho_{{\rm l},\infty}({\boldsymbol{x}},{\boldsymbol{y}})=\lim_{t\to\infty}\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}},t)$ (10)
We remove the subscript when it is clear that we are referring to the stationary density.
The calculation of $\rho_{\rm l}$ is greatly simplified if the PDE equation (6) can be solved using separation of variables, which means that the stationary solution of the form $\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}})=\nu({\boldsymbol{x}})w({\boldsymbol{y}})$. In this case the $\boldsymbol{x}$}-factor $\nu({\boldsymbol{x}})$ lies in the kernel of $\mathcal{L}$. Intuitively, we expect such a factorization to exist when $\beta$ and $h$ factorize, but as we explain below, this is not sufficient.
@@ 2.4 Population distribution
The expectation here is taken over different realizations of the population tree, which we call the population ensemble. The number density is un-normalized, but it has a very similar structure to $\rho_{\rm l}$. This is because, as long as $y_{c}>y_{b}$, the number of cells in the rectangle ${\boldsymbol{x}},{\boldsymbol{x}}+d{\boldsymbol{x}})\times[{\boldsymbol{y}},{\boldsymbol{y}}+d{\boldsymbol{y}})$ is not influenced by the division events. Meanwhile, the flux of cells in and out of this region due to the evolution of ${\boldsymbol{X}}_{t}$, the accumulation of biomass, and the division events are all unaffected by the fact that we are dealing with a growing population [53]. Therefore $n({\boldsymbol{x}},{\boldsymbol{y}},t)$ obeys equation (6). The growth only manifests at the boundary where, instead of producing one cell, a division event produces two. As a result, the boundary conditions for $n$ are
@@ 3.1 Decoupling
We will state our main result concerning decoupling. In words: under biologically reasonable factorization and independence assumptions, the long‑time joint phenotype distribution splits into the product of a size‑phenotype distribution and a growth‑phenotype distribution,
@@
where $\rho$ is either the lineage or population distribution. The precise statement of this result is given in Theorem 1 below. We impose two structural conditions:
@@
$h({\boldsymbol{x}},y_{c}|{\boldsymbol{x}}^{\prime},y_{c}^{\prime})=r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})u(y_{c}|y_{c}^{\prime})$
Condition (Condition 1) states that the division rate is scaled by the growth rate. As we will show below, this implies cells have no intrinsic age other than their size. Condition (Condition 3) asserts that growth and size are independently inherited between newborn cells. Biologically, this means that noise in the partitioning of $\boldsymbol{X}$} is independent of noise in size partitioning. Note that compared to Condition (Condition 3), Condition (Condition 1) is much stronger – not only must $\beta$ factor, but the $\boldsymbol{x}$} contribution must be exactly the growth rate.
Together, Conditions (Condition 1) and (Condition 3) imply that the integral on the right-hand side of the boundary condition (equation (8)) can be factorized into a product of functions depending separately on $\boldsymbol{x}$} and $\boldsymbol{y}$}, provided that $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ is of the separable form $\rho({\boldsymbol{x}},{\boldsymbol{y}})=\nu({\boldsymbol{x}})w({\boldsymbol{y}})$. Both assumptions are necessary for this factorization because $\beta$ appears in the boundary conditions.
However, even when both conditions hold, $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ does not necessarily separate. This is because, in our model, cell-cycle duration is determined by size. Thus, the time over which $\boldsymbol{X}$} evolves between divisions depends on size, and consequently, the distribution of $\boldsymbol{X}$} at division does too. Only when the distribution of $\boldsymbol{X}$} is invariant under the boundary conditions can we expect to have some form of decoupling. We distinguish between strong and weak forms:
@@
  • Weak Decoupling in lineage (WDl): The iteration
@@
    and $\nu=\lim_{k\to\infty}\nu_{k}\in{\rm ker}\mathcal{L}$.
In the case of SD, equation (15) holds for both $\rho_{\rm l}$ and $\rho_{\rm p}$. The size distributions $w_{\rm l}$ and $w_{\rm p}$ obey the same equations as the size-only model, while $\nu_{\rm l}\in{\rm ker}(\mathcal{L})$ and the population growth rate is determined by the eigenvalue problem, equation (16). This is, in fact, the equation for the population density and asymptotic growth rate we would arrive at if we began with a model where there was no size regulation and cells divided at a rate $\lambda$. While it is a strong assumption that $\boldsymbol{X}$} is unaffected by division events, recent analysis of high-resolution biomass data from cancer cells has shown this to be a good model of their growth rate fluctuations [27].
For WDl (resp. WDp), equation (15) holds only for the lineage (resp. population) distribution. Basically, when $\boldsymbol{X}$} is not continuous across division, we need the evolution of its probability density to be preserved in order to have some form of decoupling. However, since this evolution differs between the population and lineage distributions by a dilution term, we cannot have that both distributions decouple in the same model (except for the trivial case of constant $\lambda$).
Note that the operator $R_{\lambda}$ is not the stochastic operator, $R_{1}$, corresponding to the integral kernel $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})$; however, it is similar to $R_{1}$ in the sense that $R_{\lambda}=\lambda^{-1}R_{1}\lambda$ (here $\lambda$ is interpreted as the multiplication operator which multiples densities on $L^{1}(\mathbb{R}^{d})$ by $\lambda({\boldsymbol{x}})$). The existence of a stable, positive fixed point to the iteration equation (17) is guaranteed when $R_{1}$ is the generator of discrete-time, irreducible, ergodic Markov chain. Moreover, if $\pi$ is probability density satisfying $R_{1}\pi=\pi$, the normalized fixed point of $\nu$ satisfying $R_{\lambda}\nu=\nu$ satisfies $\nu({\boldsymbol{x}})=\pi({\boldsymbol{x}})/\lambda({\boldsymbol{x}})/\mathbb{E}_{\pi}[1/\lambda({\boldsymbol{x}})]$. This mapping between densities is known as the Esscher Transform of $\pi$ (see [12]) and is usually written in terms of an energy function $E({\boldsymbol{x}})=\ln\lambda({\boldsymbol{x}})$. This transformation is also related to the Feynman–Kac formula, which interestingly emerges separately from the population statistics, as we describe below.

### v2

@@ Abstract
We study a size-structured population model in which individual cells grow at a rate determined by a fluctuating internal variable (e.g., gene expression levels). Many previous models of phenotypically heterogeneous populations can be viewed as special cases of this model, and it has previously been observed that the internal variable decouples from cell size under certain conditions. In this work, we generalize these results and connect them to Feynman-Kac formula, which yields relationships between the lineage dynamics and population distribution in branching processes. To this end, we derive conditions for decoupling, both in the lineage and population ensemble. When decoupling occurs in both ensembles, the size dynamics can be transformed, via a random time change, into a growth-homogeneous process, and expectations can be evaluated through an exponential tilting procedure that follows from the Feynman–Kac formula. We further characterize weaker, ensemble-specific forms of decoupling that hold in either the lineage or the population ensemble, but not both. We provide a more general interpretation of tilted expectations in terms of the mass-weighted phenotype distribution.
Contents
@@ 1 Introduction
There is a vast literature on these questions [35, 38, 25, 30, 45, 51, 71, 22, 29, 27, 33, 53, 31, 4, 34, 21, 23, 14, 32, 36, 64, 63], which can be traced back to foundational work by Euler – which is summarized in the historical survey [3]. In the mathematics literature, work on these topics falls under the umbrella of structured population models. Sharpe and Lotka were the first to study branching models with age dependence [59]. Powell later explicitly connected these models with experimental measurements of microbial population growth [54]. Many others have explored mathematical questions, such as existence and uniqueness, primarily using tools from semigroup theory [16, 68, 69]. The importance of size regulation – the process by which feedback from size to division dynamics maintains stable size distributions [61] – was only recently appreciated due to work by Lin and Amir [44]. This has motivated considerable efforts in the biophysics community [50, 71, 31, 23, 67, 55]. These studies mostly focus separately on the dynamics of a fluctuating phenotype or on the mechanisms of cell-size regulation.
In this paper, we study a model of size-regulated growth in which the per unit biomass growth rate, denoted $\lambda$, is determined by internal variables ${\boldsymbol{X}}\in\mathbb{R}^{d}$ (e.g., expression levels) that evolve independently of cell size between divisions, and may be affected at division. We further assume that $\boldsymbol{X}$} may affect the division rate, denoted $\beta$; see Figure 1 A and B. Within this model, we explore (1) the structure of the stationary distribution of phenotypes in both the lineage and population ensemble, and (2) representations of population statistics in terms of lineage statistics, which are based on the Feynman–Kac formula. Previous work on Feynman–Kac representations has focused on models lacking the size variable [48], but a basic fact is that the division rate $\beta$ cannot be independent of size. This is most easily seen in a model where $\lambda$ is fixed (see Section 4.1). In the absence of any feedback from size to the division rate, the log size will undergo a random walk, and therefore the variance in sizes will diverge [1]. Meanwhile, the dependence of $\lambda$ and $\beta$ on $\boldsymbol{X}$} generates a potentially complex correlation structure between size and the phenotype $\boldsymbol{X}$} within a population. This makes the problem of determining the distribution of $\boldsymbol{X}$} difficult, because it requires marginalization over the size distribution.
@@
In the model studied in [26], division is prescribed by a random threshold that may depend on the previous cell sizes. Our motivation was to understand more generally the relationship between the size and phenotype distributions. This is achieved by building upon the adder-sizer model in references [70, 2] by combining it with the fluctuating growth rate model studied in [26, 41, 62, 27].
@@
Figure 1: (A) A diagram of the model. A cell has size ($Y$) and growth phenotypes ($X$). Division occurs at a rate that depends on both. We have omitted the explicit dependence on the initial size in this figure for simplicity. (B) A simulation of our model in the case when $X$ is an OU process, showing the evolution of $X$ (blue) and size $Y$ (red). (C) A diagram of a growing population showing the distinction between the lineage and population distributions.
@@ 1.1 Organization of this paper
$\rho_{\ell},\,\rho_{\rm p}$ Lineage and population joint densities; under decoupling, $\rho({\boldsymbol{x}},{\boldsymbol{y}})\!=\!\nu({\boldsymbol{x}})w({\boldsymbol{y}})$
$\nu_{\ell},\,\nu_{\rm p}$ Lineage and population growth‑phenotype densities
$w_{\ell},\,w_{\rm p}$ Lineage and population size‑phenotype densities
$\mathbb{E}_{\ell}$, $\mathbb{E}_{\rm p}$ expectations with respect to $\rho_{\ell}$ and $\rho_{\rm p}$
@@ 2 Size and phenotype structured population model
Here, we describe a class of single cell growth models in which a cell’s phenotype at any instant can be decomposed into growth phenotype ${\boldsymbol{X}}(t)\in\mathbb{R}^{d}$ and size phenotype ${\boldsymbol{Y}}(t)=(Y_{c}(t),Y_{b}(t))\in\mathbb{R}^{2}$. The growth phenotype, which could represent the concentration of growth-limiting cellular components such as ribosomes or proteins involved in antibiotic resistance, evolves independently of size between cell divisions according to some Markov process as described below. The size phenotype consists of the normalized (log) current cell size and size at the most recent division. Specifically, if $M_{b}(t)$ is the size at birth of a cell at time $t$, we set ${\boldsymbol{Y}}(t)=(Y_{c}(t),Y_{b}(t))$ where $Y_{c}=\ln(M(t)/\mathbb{E}[M_{b}])$ and $Y_{b}(t)=\ln(M_{b}(t)/\mathbb{E}[M_{b}])$, and $\mathbb{E}[M_{b}]$ is the mean steady-state birth size. Meanwhile, cells accumulate size (e.g. mass) $M(t)=e^{Y_{c}(t)}\mathbb{E}[M_{b}]$ exponentially at a rate $\lambda({\boldsymbol{X}}(t))$; i.e.,
$Y_{c}^{\prime}(t)=\lambda({\boldsymbol{X}}(t)),$ (1)
and divide at a rate $\beta({\boldsymbol{X}}(t),{\boldsymbol{Y}}(t))$ yielding two newborn cells whose phenotypes are sampled from a distribution $h({\boldsymbol{x}},y_{c}|{\boldsymbol{x}}^{\prime},y_{c})$.
Note that the initial size is expressed as a function of time because it changes during cell division but remains constant within a cell cycle. The motivation for keeping track of both the current size and size at birth is that many cells exhibit size-regulation mechanisms in which the growth increment between divisions depends on the initial size. To accommodate such regulatory strategies within our modeling framework, it is necessary to include both the current cell size and initial size.
@@ 2.1 Growth phenotype dynamics
propagate expectations; that is, $g^{f}({\boldsymbol{x}},t)=\mathbb{E}[f(\tilde{{\boldsymbol{X}}}(t))|\tilde{{\boldsymbol{X}}}(0)={\boldsymbol{x}}]$ with $g^{f}({\boldsymbol{x}},0)=f({\boldsymbol{x}})$. We use the convention that $\mathcal{L}$ (without the adjoint notation) propagates the probability density because this will appear more often in our presentation. As explained later, we will allow the possibility that $\mathcal{L}=0$, in which case $\tilde{{\boldsymbol{X}}}(t)=\tilde{{\boldsymbol{X}}}(0)$.
@@ 2.2 Size and division dynamics
where $t_{k}$ is the $k$th division event, $({\boldsymbol{X}}(t_{k}^{-}),Y_{c}(t_{k}^{-}))$ are the growth and size phenotypes immediately before division, and ${\boldsymbol{x}},{\boldsymbol{x}}+d{\boldsymbol{x}})$ is shorthand for the rectangle $x_{1},x_{1}+dx_{1})\times\cdots\times[x_{d},x_{d}+dx_{d})$.
@@
  • A linear operator $\mathcal{L}$ that is the generator of a continuous time Markov process (including the trivial case $\mathcal{L}=0$).
@@
  • A transition kernel $h:(\mathbb{R}^{d}\times\mathbb{R})\times(\mathbb{R}^{d}\times\mathbb{R})\to\mathbb{R}_{\geq 0}$ gives the phenotypes of a cell following a division event. $h(\cdot,\cdot|{\boldsymbol{x}},y_{c})$ is a probability density that respects the conservation of mass at division– that is, the sum of the two newborn daughter cell sizes must equal the mother cell size immediately before division. Said another way, if $y_{c}^{\prime}$ is the log of the size of the mother cell before division and $y_{c}$ is the size of a daughter cell, then $e^{y_{c}^{\prime}}-e^{y_{c}}$ (the mass of the other daughter cell) is equal in distribution to $e^{y_{c}}$.
@@ 2.3 Lineage distribution
The lineage distribution is obtained when we consider the distribution of phenotypes over a sequence of cells, where at each division, one of the daughter cells is discarded. The evolution of the joint density $\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})$ of growth and size phenotypes, $({\boldsymbol{X}}_{t},{\boldsymbol{Y}}_{t})=({\boldsymbol{x}},{\boldsymbol{y}})$ along a lineage obeys the PDE
$\begin{split}\partial_{t}\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}},t)&=\mathcal{A}\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}},t)\end{split}$ (7)
@@
The advection term comes from equation (1).
The boundary condition at $y_{c}=y_{b}$ is obtained by examining the flux of probability mass in the $y_{c}$ direction at zero added size. Newborn cells have $y_{c}=y_{b}$, and because cells do not divide at the instant of birth, there is no loss of mass through this boundary. The only inflow comes from division events in the interior with $y_{c}^{\prime}>y_{b}$ while the deterministic increase of added size generates an advective flux through the boundary equal to
$\lambda({\boldsymbol{x}})\,\rho_{\mathrm{l}}^{b}({\boldsymbol{x}},y_{b},t),$ (9)
where $\rho_{\mathrm{l}}^{b}({\boldsymbol{x}},y_{b},t)$ denotes the trace of the density at $y_{c}=y_{b}$. A mother cell with phenotype $({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime})$ divides at rate $\beta({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime})$ and produces a newborn daughter with phenotype $({\boldsymbol{x}},y_{b})$ with density $h({\boldsymbol{x}},y_{b}\mid{\boldsymbol{x}}^{\prime},y_{c}^{\prime})$. Integrating over all possible mother states gives the total rate at which newborns of type $({\boldsymbol{x}},y_{b})$ are created. Equating this influx with the outgoing advective flux yields
$\begin{split}\lambda({\boldsymbol{x}})\,\rho_{\mathrm{l}}^{b}({\boldsymbol{x}},y_{b},t)=\int_{y_{b}}^{\infty}\!\int_{-\infty}^{y_{c}^{\prime}}\!\int_{\mathbb{R}^{d}}h({\boldsymbol{x}},y_{b}\mid{\boldsymbol{x}}^{\prime},y_{c}^{\prime})\,\beta({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime})\,\rho_{\mathrm{l}}({\boldsymbol{x}}^{\prime},y_{c}^{\prime},y_{b}^{\prime},t)\,d{\boldsymbol{x}}^{\prime}\,dy_{b}^{\prime}\,dy_{c}^{\prime}.\end{split}$ (10)
Throughout, the superscript $b$ is used to denote the distribution of newborn cells.
We are interested in the steady-state distributions
$\rho_{{\ell},\infty}({\boldsymbol{x}},{\boldsymbol{y}})=\lim_{t\to\infty}\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}},t)$ (11)
We remove the subscript when it is clear that we are referring to the stationary density. The calculation of $\rho_{\ell}$ is greatly simplified if the PDE equation (7) can be solved using separation of variables, which means that the stationary solution of the form $\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})=\nu({\boldsymbol{x}})w({\boldsymbol{y}})$. In this case the $\boldsymbol{x}$}-factor $\nu({\boldsymbol{x}})$ lies in the kernel of $\mathcal{L}$. Intuitively, we expect such a factorization to exist when $\beta$ and $h$ factorize, but as we explain below, this is not sufficient.
@@ 2.4 Population distribution
The expectation here is taken over different realizations of the population tree, which we call the population ensemble. The number density is un-normalized, but it has a very similar structure to $\rho_{\ell}$. This is because, as long as $y_{c}>y_{b}$, the number of cells in the rectangle ${\boldsymbol{x}},{\boldsymbol{x}}+d{\boldsymbol{x}})\times[{\boldsymbol{y}},{\boldsymbol{y}}+d{\boldsymbol{y}})$ is not influenced by the division events. Meanwhile, the flux of cells in and out of this region due to the evolution of ${\boldsymbol{X}}_{t}$, the accumulation of biomass, and the division events are all unaffected by the fact that we are dealing with a growing population [70]. Therefore $n({\boldsymbol{x}},{\boldsymbol{y}},t)$ obeys equation (7). The growth only manifests at the boundary where, instead of producing one cell, a division event produces two. As a result, the boundary conditions for $n$ are
@@ 3.1 Decoupling
We will state our main result concerning decoupling. In words: under biologically reasonable factorization and independence conditions, the long‑time joint phenotype distribution splits into the product of a size‑phenotype distribution and a growth‑phenotype distribution,
@@
where $\rho$ is either the lineage or population distribution. As with $\rho$ will use the subscripts $\rm p$} and $\ell$} on $\nu$ and $w$ to indicate the population and lineage densities. The precise statement of this result is given in Theorem 1 below. We impose two structural conditions, which will be referenced throughout.
@@
$h({\boldsymbol{x}},y_{b}|{\boldsymbol{x}}^{\prime},y_{c}^{\prime})=r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})u(y_{b}|y_{c}^{\prime})$
Condition (Condition 1) is equivalent to assuming that the division hazard per unit of (log-)size accumulation depends only on the size variables: since $dY_{c}/dt=\lambda(X(t))$, the “hazard per unit added log-size” is $\beta/\lambda=\phi$. Thus, fluctuations in $\lambda$ modulate the speed of progression through the size-controlled division process rather than introducing an independent “chronological age” clock. This is consistent with standard size-control modeling (adder/sizer-type regulation) [31, 1] and it allows the growth process to be transformed to a process with constant growth rate via a random time-change $(T(t)=\int_{0}^{t}\lambda(X(s)),ds)$ (see Section 5).
Condition (Condition 3) corresponds to independent inheritance noise in the internal phenotype versus size partitioning noise, which may be reasonable in some settings but not universal. Section 6 discusses how we can still obtain lineage representations of the population distribution when this condition fails.
Note that compared to Condition (Condition 3), Condition (Condition 1) is much stronger; not only must $\beta$ factor, but the $\boldsymbol{x}$} contribution must be exactly the growth rate.
Together, Conditions (Condition 1) and (Condition 3) imply that the integral on the right-hand side of the boundary condition (equation (10)) can be factorized into a product of functions that depend separately on $\boldsymbol{x}$} and $\boldsymbol{y}$}, provided that $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ is of the separable form $\rho({\boldsymbol{x}},{\boldsymbol{y}})=\nu({\boldsymbol{x}})w({\boldsymbol{y}})$. Both conditions are necessary for this factorization because $\beta$ appears in the boundary conditions. However, even when both conditions hold, $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ does not necessarily separate. This is because cell-cycle duration is determined by size. Thus, the time over which $\boldsymbol{X}$} evolves between divisions depends on size, and consequently, the distribution of $\boldsymbol{X}$} at division does too. Only when the distribution of $\boldsymbol{X}$} is invariant under the boundary conditions can we expect to have some form of decoupling. We distinguish between strong and weak forms:
@@
  • Weak Decoupling (WD): The iteration
@@
    and $\nu=\lim_{k\to\infty}\nu_{k}\in{\rm ker}\mathcal{L}$. The most biologically relevant case is when $\mathcal{L}=0$, in which case this model becomes the random growth rate model discussed in Section 4.2.
In the case of SD, equation (16) holds for both $\rho_{\ell}$ and $\rho_{\rm p}$. The size distributions $w_{\ell}$ and $w_{\rm p}$ obey the same equations as the size-only model, while $\nu_{\ell}\in{\rm ker}(\mathcal{L})$ and the population growth rate is determined by the eigenvalue problem, equation (17). This is, in fact, the equation for the population density and asymptotic growth rate we would arrive at if we began with a model where there was no size regulation and cells divided at a rate $\lambda$. While it is a strong assumption that $\boldsymbol{X}$} is unaffected by division events, recent analysis of high-resolution biomass data from cancer cells has shown this to be a good model of their growth rate fluctuations [40]. Note that the case of a constant growth rate $\lambda({\boldsymbol{x}})=\lambda_{0}$ is also covered by SD, since Condition (Condition 1) automatically holds here.
For WD, equation (16) holds only for the lineage distribution. When $\boldsymbol{X}$} is not continuous across division, we need the evolution of its probability density to be preserved in order to achieve some form of decoupling. However, since this evolution differs between the population and lineage distributions by a dilution term, we cannot have both distributions decouple in the same model (except for the trivial case of constant $\lambda$).
Note that the operator $R_{\lambda}$ is not the stochastic operator, $R_{1}$, corresponding to the integral kernel $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})$; however, it is similar to $R_{1}$ in the sense that $R_{\lambda}=\lambda^{-1}R_{1}\lambda$ (here $\lambda$ is interpreted as the multiplication operator that multiplies densities on $L^{1}(\mathbb{R}^{d})$ by $\lambda({\boldsymbol{x}})$). The existence of a stable, positive fixed point to the iteration equation (18) is guaranteed when $R_{1}$ is the generator of a discrete-time, irreducible, ergodic Markov chain. Moreover, if $\pi$ is probability density satisfying $R_{1}\pi=\pi$, the normalized fixed point of $\nu$ satisfying $R_{\lambda}\nu=\nu$ satisfies $\nu({\boldsymbol{x}})=\pi({\boldsymbol{x}})/\lambda({\boldsymbol{x}})/\mathbb{E}_{\pi}[1/\lambda({\boldsymbol{x}})]$. This mapping between densities is known as the Esscher Transform of $\pi$ (see [15]) and is usually written in terms of an energy function $E({\boldsymbol{x}})=\ln\lambda({\boldsymbol{x}})$. This transformation is also related to the Feynman–Kac formula, which interestingly emerges separately from the population statistics, as we describe below.
Finally, we note that a third case, which we refer to as Weak Decoupling in the population (WDp), may be considered. Here, decoupling occurs only in the population. However, because this does not have any natural biological meaning, we reserve a discussion to the technical details in Section 7.